Express each number as a product of its prime factors:
5005

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

5005
By Taking the LCM of 5005, we will get the product of its prime factor.
Hence,5005 = 5 × 7 × 11 × 13 × 1 = 5 × 7 × 11 × 13

​Thanks